Use Substitution to solve each system of equations. If the system does not have exactly one solution, state if it has no solution or infinitely many solutions. x-5y=36 and 2x+y=-16
Answer by eperette(173)

x - 5y = 36
2x + y = -16
before we can use the substitution method, we should have an x= or y= equation....so
x - 5y = 36
x-5y+5y=36+5y
x+0 = 36 + 5y
x= 5y + 36
Now we will substitute the right part of this equation into the x value in the second equation
2x + y = -16
2(5y + 36) + y =-16
2(5y) + 2(36) + y = -16
10y + 72 + y = -16
11y + 72 = -16
11y+72-72=-16-72
11y+0=-88
11y/11=-88/11
y=-8
Now to find x, we will substitute y=-8 into first equation
x = 5y + 36
x = 5(-8) + 36
x = -40 + 36
x = -4
So my answer is (-4, -8)...
Lets check:
x-5y=36
(-4)-5(-8)
-4 + 40
36 right
AND
2x+y=-16
2(-4)+(-8)
-8 -8
-16 right